OpenStudy (anonymous):

If the sin of angle x is 4/5 , and the triangle was dilated to be two times as big as the original, what would be the value of the sin of x for the dilated triangle? Hint: Use the slash symbol ( / ) to represent the fraction bar and enter the fraction with no spaces.

OpenStudy (anonymous):

Sin of the angle is unchanged regardless of the size; it is a ratio related only to the angle, not the size of the triangle.

Directrix (directrix):

The dilated triangle is similar to the original one. So, the ratio of corresponding sides will remain the same. By definition of similar triangles, corresponding sides are in proportion.

OpenStudy (anonymous):

if it remians the same would it still be 4/5?

Directrix (directrix):

Yes.

Directrix (directrix):

When the triangle is dilated, the 4 side jumps up to 8 and the 5 side jumps up to 10. Yet, the ration of 8/10 is still equal to 4/5.

Directrix (directrix):

At the core of our discussion is this attached theorem which ensures that the two triangles are similar.

OpenStudy (anonymous):

do you know the answer to this one??c: If the sin 30° is 1/2 , then the cos ____° = _____. a. 60° ; 1/2 b. 30° ; \[\frac{ \sqrt{2} }{ 2 }\] c. 60° ; \[\frac{ \sqrt{3} }{ 2 }\] d. 30° ; 1

Directrix (directrix):

The trig functions Sine and Cosine are co-functions. Do you know what that means within the context of Trigonometry?

OpenStudy (anonymous):

yes, but im checking my anwsr on this one

Directrix (directrix):

What do you think is the answer to just this part: sin 30° = cos ____° ? http://planetandepoch.com/?p=1488

Directrix (directrix):

Study Guide

Directrix (directrix):

@lowcard2 What do you think is the answer to just this part: sin 30° = cos ____° ?

OpenStudy (anonymous):

i put c

Directrix (directrix):

If I may ask, how did you rule out this option: a. 60° ; 1/2 @lowcard2

OpenStudy (anonymous):

my friend showed me steps and the answer

OpenStudy (anonymous):

what is the correct answr then?

Directrix (directrix):

I think I misread the question. Hold on a second.

Directrix (directrix):

> i put c That is not correct.

OpenStudy (anonymous):

what is it then?

Directrix (directrix):

Sine and Cosine are cofuntions where "co" is for complementary angles. So, sin(30) = cos(90 - 30) = cos(60) So, if sin(30) = 1/2, the cos(60) = the same which is 1/2. @lowcard2 Does that make sense? It is all about the cofunctions.

OpenStudy (anonymous):

yes it does thank you

Directrix (directrix):

@lowcard2 So that we don't get our wires crossed, which option is the answer to this question.

OpenStudy (anonymous):

i put a

Directrix (directrix):

I agree. When you have time, take a look at the study attachments. I think they will help clarify your thinking. Trig is tricky.